OpenStudy (ujjwal):

In an Argand plane, the area in square units of the triangle formed by the points (1+i), (1-i), 2i is:

OpenStudy (ujjwal):

I think the answer should be i square units. But my book says its 1/2 square units.

OpenStudy (asnaseer):

@ujjwal - do you really mean that you think the answer is "i square units", or did you mean "one square unit"?

OpenStudy (ujjwal):

I mean i square units..

OpenStudy (asnaseer):

how?

OpenStudy (ujjwal):

 Area= 1/2 \(\times\) base \(\times\) height= 1/2 \(\times\) 2i \(\times\) 1=i